OpenStudy (anonymous):

Which is the equation of the line passing through (0, 4) and (8, 0)? 1.y= -1/2x + 4 2.y= -1/2x + 8 3.y=-2x + 4 4.y=-2x + 9

OpenStudy (e.mccormick):

Know how to find the slope using any two points?

OpenStudy (anonymous):

well the slopes for 3. and 4. would be -2/1 right?

OpenStudy (e.mccormick):

Yes. But I am saying, start with (0, 4) and (8, 0) and find the slope. Know how to do that?

OpenStudy (e.mccormick):

\(\dfrac{\Delta y}{\Delta x}\) AKA: \(\dfrac{y_2-y_1}{x_2-x_1}\)

OpenStudy (anonymous):

so 4/0 and 0/8 right?

OpenStudy (e.mccormick):

No. If (0, 4) is \((x_1, y_1)\) and (8, 0) is \((x_2, y_2)\) what do you get when you substitude them into \(\dfrac{y_2-y_1}{x_2-x_1}\)? Ypu can copy and paste `\(\dfrac{y_2-y_1}{x_2-x_1}\)` if you want it formatted nice like what I type up. Or use (y_2-y_1)/(x_2-x_1) and replace them with the right numbers. Then you simplify.

OpenStudy (anonymous):

(0-4)/(8-0)

OpenStudy (anonymous):

-4/8? or 4/8

OpenStudy (e.mccormick):

Right. Which becomes -4/8 and simplifies a bit more than that.

OpenStudy (anonymous):

-1/2

OpenStudy (anonymous):

so its either 1 or 2 now

OpenStudy (e.mccormick):

Yes. So the slope, which is sometimes called m, is -1/2. OK, so next, you need to find the intercept. \(y=mx+b\) You know m, you need b. You have been given two points and found m. With that, do you think you can find b?

OpenStudy (anonymous):

i forgot how to find b again haha

OpenStudy (e.mccormick):

Pick one of the two points. Use the x and y from that point, as well as the m you just found, and plug them into \(y=mx+b\). See, \(y=mx+b\) has 4 variables. But you have 3 of the 4. Suddently you are able to solve for the remainging 1 when you plug in the other 3.

OpenStudy (e.mccormick):

This is very typical in math. You use what is known to find what is unknown. You were given two points. You need \(y=mx+b\) form, so you need m and b. The two points let you find m. Then one point and m let you find b. Once you know m and b, then you can see the proper \(y=mx+b\) form that answers this question. It is just a set of steps. Each one is a small step. They build up as you do them. Eventually they make a solution.

OpenStudy (anonymous):

hm

OpenStudy (anonymous):

y=-1/2x +b im still confused on how to find b

OpenStudy (e.mccormick):

Personally I like the first point because x is easy in that one.

OpenStudy (anonymous):

so its y=-1/2x +4

OpenStudy (e.mccormick):

Yep!